Table 1            Estimated Total Returns                                                                                                 
State of the Economy Probability T-Bond SETX Golden S&P 500
Recession 5% 5% -19% 20% -14%
Below Average 15% 5% 2% 13% 3%
Average 45% 5% 9% 10% 11%
Above Average 25% 5% 34% 5% 22%
Boom 10% 5% 25% -5% 33%

Golden expects to have $1.45 in earnings next year. Golden had $1.30 in earnings 3 years ago. What should Golden sell for?
Answer #1

Growth Rate = (1.45/1.30)1/4 - 1 = 2.77%

Expected Rate = 0.05(0.20) + 0.15(0.13) + 0.45(0.10) + 0.25(0.05) + 0.10(-0.05)

Expected Rate = 8.20%

Using DDM Model,

Stock Price = 1.45/(0.082 - 0.0277)

Stock Price = $26.70